# A multi-messenger search for exotic field emission with a global magnetometer network

### v1

@@ V.2 Limits
With no significant events detected, we determine a threshold $\lambda^{th}_{n}$ for each box $n$ corresponding to an expected value of one false positive over the dataset, i.e. $\text{FPR}_{n}(\lambda^{th}_{n})=1/N^{fg}_{tot}$ [Eq. (32)]. These are then translated into a 90%-confidence upper limit on $\mathcal{B}_{\text{p}}$ with the FC belts constructed for each box which are displayed in Fig. 10.
@@
The 90%-confidence exclusion limits on the energy scale $f^{l(q)}$ [Eqs. (15)-(16)] are determined with the 90%-confidence upper limit on $\mathcal{B}_{\text{p}}$ and the center delay time and frequency of each box. Figure 11 displays the exclusion limits calculated for $\tau_{0}=1$ s and $\Delta E=1\,M_{\odot}c^{2}$.
The exclusion limits can be calculated for other values of $\Delta E$ by simply multiplying our presented results by $\sqrt{\Delta E/M_{\odot}c^{2}}$. The exclusion limits depend on $\tau_{0}$ through Eq. (6), which depends on the delay time and frequency, and can be scaled to other values of $\tau_{0}$ by recalculating $\tau$ at the new value for each box. For all boxes, there is a limit on the initial burst duration, $\tau_{0}>1/\omega_{0}$, for the emitted energy spectra to be sufficiently sharp in frequency space Dailey2021 .
Finally, we assume the ELF pulse and GW/EM trigger are emitted at the same time. If an alternative production theory postulates the time of emission of the ELFs and GW/EM differ then an appropriate offset should be added to the delay time.
@@ V.3 Projected sensitivity to other events
The GNOME collaboration is upgrading station sensors to co-magnetometers for the “Advanced GNOME” Afach2023 that boast orders of magnitude higher sensitivity to ELF signals PhysRevResearch.6.013339_comag . Figure 12 displays the projected sensitivity of this analysis to the ELF energy scales across a range of astrophysical distances. The curves are calculated for the approximate network sensitivity in this work and the projected network sensitivity of Advanced GNOME. The associated energy scales for linear and quadratic ELF interactions [Eqs. (15)-(16)] are evaluated at $\tau_{0}=1$ s, $\omega_{0}/2\pi=1$ Hz, and $\Delta E=1\,M_{\odot}c^{2}$ over a range of astrophysical distances. GNOME could search parameter space beyond the astrophysical limits with the next generation of co-magnetometer sensors.
@@
Figure 12: Network sensitivity of this work and projected sensitivity of Advanced GNOME. The curves are Eqs. (15) and (16) calculated with $\tau_{0}=1$ s, $\omega_{0}/2\pi=1$ Hz, and $\Delta E=1\,M_{\odot}c^{2}$ for different values of $\mathcal{B}_{\text{p}}$ and $\delta t$. The astrophysical limits on $f_{l}$ and $f_{q}$ are $2\times 10^{8}$ GeV Chang2018 and $1\times 10^{4}$ GeV pospelov2013detecting , respectively. The vertical dotted line indicates the distance to the BBH merger S200311b probed in this paper.
VI Conclusions
@@
We have developed an algorithm to search GNOME data for an exotic low-mass field burst copropagating with gravitational or electromagnetic signals from large-energy astrophysical events. We applied this analysis to data coincident with the BBH merger S200311bg detected by LIGO/Virgo. No significant events were detected and we consequently placed the first lab-based constraints on combinations of ELF production and coupling parameters, namely the energy scales of the interaction between the ELF and proton spins, the energy released from the astrophysical event in the form of an ELF burst, and the initial ELF-burst duration.
@@
This analysis will be applied to more GNOME datasets that are coincident with other gravitational-wave events detected by LIGO/Virgo and dozens of fast radio bursts. This will expand the mass range explored and closer events will allow higher energy scales to be probed. The collaboration is upgrading to self-compensating noble-gas-alkali-metal comagnetometers for implementation in the Advanced GNOME experiment. These improved sensors will be able to probe the coupling of exotic fields to protons and neutrons with higher sensitivity than our current sensors which can only probe the coupling to protons. Additionally, the response to magnetic and non-magnetic signals in the comagnetometers can be discriminated allowing for more efficient suppression of false positives. Furthermore, this analysis technique could be generalized to an all-sky search to hunt for potential signals without a known trigger.
@@ Acknowledgements
This work was supported by the U.S. National Science Foundation under grants PHY-1707875, PHY-1707803, PHY-2110370, PHY-2110388, and PHY-2207546; by the German Research Foundation (DFG) under grant no. 439720477 and within the German Excellence Strategy (Project ID 39083149); by work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology)”; ZDG acknowledges institutional funding provided by the Institute of Physics Belgrade through a grant by the Ministry of Science, Technological Development and Innovations of the Republic of Serbia. The work of SP is supported by the National Science Centre of Poland within the Opus program (grant 2019/34/E/ST2/00440) and GL acknowledges the support of the Excellence Initiative – Research University of the Jagiellonian University in Kraków.

### v2

@@ V.2 Limits
With no significant events detected, we place exclusion limits on the energy scales of the linear and quadratic ELF $\phi$-spin interaction. We determine a threshold $\lambda^{th}_{n}$ for each box $n$ corresponding to an expected value of one false positive over the dataset, i.e. $\text{FPR}_{n}(\lambda^{th}_{n})=1/N^{fg}_{tot}$ [Eq. (32)]. These are then translated into a 90%-confidence upper limit on $\mathcal{B}_{\text{p}}$ with the FC belts constructed for each box which are displayed in Fig. 10.
@@
The 90%-confidence exclusion limits on the energy scale $f^{l(q)}$ [Eqs. (15)-(16)] are determined with the 90%-confidence upper limit on $\mathcal{B}_{\text{p}}$ and the center delay time and frequency of each box. Figure 11 displays the exclusion limits calculated for $\tau_{0}=1$ s and $\Delta E=1\,M_{\odot}c^{2}$. The exclusion limits are the first lab-based constraints placed on the energy scales of the ELF $\phi$-spin interaction. From the energy release, initial burst duration, and distance of a given astrophysical source, we can place a corresponding limit on the ELF parameter space in the case of a non-detection.
The exclusion limits can be calculated for other values of $\Delta E$ by simply multiplying our presented results by $\sqrt{\Delta E/M_{\odot}c^{2}}$. The exclusion limits depend on $\tau_{0}$ through Eq. (6), which depends on the delay time and frequency, and can be scaled to other values of $\tau_{0}$ by recalculating $\tau$ at the new value for each box. For all boxes, there is a limit on the initial burst duration, $\tau_{0}>1/\omega_{0}$, for the emitted energy spectra to be sufficiently sharp in frequency space Dailey2021 ; arakawa2025multimessenger .
Finally, we assume the ELF pulse and GW/EM trigger are emitted at the same time, for example due to a bosenova/relativistic axion burst takahashi2024self ; aurrekoetxea2024self ; eby2022probing ; arakawa2024detection . If an alternative production theory postulates the time of emission of the ELFs and GW/EM differ then a simple timing offset can be added to the analysis.
These limits allow us to explore new territory in both particle physics and astrophysics. From the perspective of BSM physics, our search probes three key aspects of exotic field models: (1) the existence of axion-like particles (ALPs) or other low-mass bosons, (2) their coupling to Standard Model fermion spins, and (3) their potential production in high-energy astrophysical events. While previous laboratory and astrophysical searches have placed bounds on ALP couplings and relic populations, the scenario of coherent bursts produced in dynamic astrophysical environments remains largely unconstrained. Our exclusion plots (Figs. 10 and 11) do not yet reach the coupling strengths excluded by other laboratory or astrophysical observations cong2024spin ; Raf99 ; antypas2022new , but they begin to constrain the combined parameter space of ELF production and coupling. These constraints provide valuable guidance for theory development in this still-nascent area arakawa2025multimessenger , including models of bosonic clouds around compact objects arvanitaki2011exploring ; baryakhtar2021black or bosenova collapse mechanisms arakawa2024detection . More broadly, ELF detection or exclusion enables physical insight into the dynamics and environments of extreme astrophysical phenomena. For example, coincident ELF signals could reveal new channels of energy release in black hole or neutron star mergers alvarez2024finalparsec , or indicate the presence of novel fields involved in supernova explosions Raf88 or fast radio bursts Iwa15 ; Tka15 . Thus, even null results, as presented here, help narrow the range of viable models and open the door to new observational probes of both fundamental physics and the astrophysical processes that produce them.
@@ V.3 Projected sensitivity to other events
The GNOME collaboration is upgrading station sensors to co-magnetometers for the “Advanced GNOME” network Afach2023 that boast orders of magnitude higher sensitivity to ELF signals PhysRevResearch.6.013339_comag . Figure 12 displays the projected sensitivity of this analysis to the ELF energy scales across a range of astrophysical distances. The curves are calculated for the approximate network sensitivity in this work and the projected network sensitivity of Advanced GNOME. The associated energy scales for linear and quadratic ELF interactions [Eqs. (15)-(16)] are evaluated at $\tau_{0}=1$ s, $\omega_{0}/2\pi=1$ Hz, and $\Delta E=1\,M_{\odot}c^{2}$ over a range of astrophysical distances. Astrophysical limits based on the rate of SN1987A cooling on $f_{l}$ and $f_{q}$ are $1.5\times 10^{9}$ GeV Lella_2024 and $1\times 10^{4}$ GeV pospelov2013detecting , respectively. GNOME could search parameter space beyond the indirect astrophysical limits with the next generation of co-magnetometer sensors.
@@
Figure 12: Network sensitivity of this work and projected sensitivity of Advanced GNOME. The curves are Eqs. (15) and (16) calculated with $\tau_{0}=1$ s, $\omega_{0}/2\pi=1$ Hz, and $\Delta E=1\,M_{\odot}c^{2}$ for different values of $\mathcal{B}_{\text{p}}$ and $\delta t$. The gray area represents the astrophysical exclusion limits on $f_{l}$ and $f_{q}$ which are $1.5\times 10^{9}$ GeV Lella_2024 and $1\times 10^{4}$ GeV pospelov2013detecting , respectively. The purple area represents the parameter space ruled out by laboratory measurements of proton-spin couplings ramsey1979tensor ; cong2024spin . The vertical dotted line indicates the distance to the BBH merger GW200311_115853 probed in this paper.
Expanding the number of detectors in the network is another route to increasing sensitivity and false positive rejection. The question of the optimal configuration of the network was explored in Ref. smiga2022assessing . Each detector makes an independent measurement of $\mathcal{B}_{\text{p}}$ based on the projection of $\mathcal{B}_{\text{p}}$ on its sensitive axis. As three stations are required to define the projection of $\mathcal{B}_{\text{p}}$ on the network, we have that to first order, the signal-to-noise ratio of the detection LRT statistic will scale as $\sqrt{N}$ for $N>3$ under the assumption of equivalent sensitivity for each detector. Corrections to this scaling – accounted for in this analysis – arise due to the particular target considered and the relative detector noise. Fig. 10 shows an effective true positive network sensitivity in $pT$ for GW200311_115853 after folding in the noise characteristics, and orientation of the 5 stations used in this search. As noted in Ref. pospelov2013detecting , increasing the number of detectors can give faster than $\sqrt{N}$ improvement for false positive rejection in the case where the background is dominated by rare non-Gaussian noise spikes.
VI Conclusion
@@
Using the Global Network of Optical Magnetometers for Exotic physics (GNOME), we have performed a multi-messenger astronomy search for ultralight, scalar BSM fields. This work considered the emission of an exotic low-mass field (ELF) burst produced during the binary black hole merger GW200311 from which gravitational waves were detected by LIGO/Virgo GCN_BBH . No significant events were detected by GNOME. We consequently place the first lab-based constraints on combinations of ELF production parameters, namely the energy scales of the interaction between the ELF and proton spins, the energy released from the astrophysical event in the form of an ELF burst, and the initial ELF-burst duration.
@@
In the future, this analysis method will be applied to existing GNOME datasets Afach2023 which overlap with other gravitational-wave events detected by LIGO/Virgo PhysRevX.13.041039 and dozens of fast radio bursts frbcollaboration2025cataloglocaluniversefast . This will expand the mass range explored and closer events will allow higher energy scales to be probed. The collaboration is further upgrading to self-compensating noble-gas-alkali-metal comagnetometers for implementation in the Advanced GNOME experiment klinger2023optimization ; PhysRevResearch.6.013339_comag . These improved sensors will be able to probe the coupling of exotic fields to protons and neutrons with higher sensitivity than our current sensors which can only probe the coupling to protons. Additionally, the response to magnetic and non-magnetic signals in the comagnetometers can be discriminated, allowing for a reduction of the false positive rate Afach2023 . As an extension of this work, we plan to generalize this technique to an all-sky search to hunt for potential signals without a known trigger. In the event of a detection, the network can truly act like a telescope with the ability to localize the signal origin. Any ELF detection would be indicative of new physics – in particular, an ELF observation would not only be a discovery of a BSM particle, but would give insight into as yet unknown production mechanisms in an extremely high energy and strongly curved spacetime regime.
The development of quantum sensors has led to a variety of measurements with exquisite sensitivities in individual laboratories Ye2024 . Using GPS, it is now straightforward for labs to accurately time stamp data – enabling new networks of quantum sensors. To analyze data from a different quantum sensor network using the algorithm described in this paper, one would primarily have to change the function that describes the network response to an ELF, e.g. Eq. (20), for these other sensors and account for differing sampling rates and bandwidths. We propose that the analysis framework presented here can serve as a template that can be used for a variety of networks searching for BSM physics. Multimessenger astronomy with quantum sensors will inform investigations into astrophysical production mechanisms of BSM fields, their propagation, and their interactions with matter – opening new windows with which to view the universe.
@@ Acknowledgements
This work was supported by the U.S. National Science Foundation under grants PHY-1707875, PHY-1707803, PHY-2110370, PHY-2110388, and PHY-2207546; by the German Research Foundation (DFG) under grant no. 439720477 and within the German Excellence Strategy (Project ID 39083149); by work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology)”; ZDG acknowledges institutional funding provided by the Institute of Physics Belgrade through a grant by the Ministry of Science, Technological Development and Innovations of the Republic of Serbia. The work of SP is supported by the National Science Centre of Poland within the Opus program (grant 2019/34/E/ST2/00440) and GL acknowledges the support of the Excellence Initiative – Research University of the Jagiellonian University in Kraków. The authors thank Yuanhong Wang for pointing out an error in our calculation in the preprint version of this article.